OpenStudy (anonymous):

A printer can print forty-five pages in 9 minutes. At that rate, how many pages can the printer print in fourteen minutes?

OpenStudy (anonymous):

Find the amount of pages the printer prints per minute and multiply that by 14.

OpenStudy (anonymous):

fifty pages fifty-nine pages seventy pages eighty pages

OpenStudy (e.cociuba):

Answer: C) 70 pages

OpenStudy (e.cociuba):

45 pages / 9 minutes = 5 * 14 minutes = 70 pages, or you can cross multiple 45 pages * 14 minutes = 630 / 9 minutes = 70 pages

OpenStudy (mathstudent55):

Use a proportion: -------------------- | pages | 45 | x | -------------------- | minutes | 9 | 14 | -------------------- 45:9 = x:14 45/9 = x/14 Cross multiply and solve for x.